# Spacetime Grand Unified Theory

### v1

@@ Abstract
The Standard Model of particle physics is derived from first principles starting from the free Dirac Lagrangian. All known fermionic particle species plus three right handed neutrinos are obtained from ideals of a $\mathcal{C}\ell_{8}$ algebra, with gauge symmetries arising as rotations of creation-annihilation operators and vacua. Triality originates both the strong force and three particle families with a mass hierarchy. Lorentz and gauge transformations are unified while avoiding the Coleman-Mandula theorem. Chirality stems unavoidably from rotations leaving the vacua invariant, with a predicted Weinberg angle of $\sin^{2}(\theta_{W})=3/8$. The theory is anomaly-free and devoid of proton decay.
@@ I Introduction
An algebraic alternative to GUTs and SUSY has been simultaneously developed over the decades, though it has received far less attention in comparison. This approach, typically centered on division algebras Gunaydin:1973a ; Gunaydin:1973b ; Leo:1996 ; Carrion:2003 ; Baez:2012 ; Burdik:2017 ; Furey:2022a ; Furey:2022b or Clifford algebras Casalbuoni:1979 ; Hestenes:1982 ; Trayling:2001 ; Doran:2003 ; Gording:2020 ; Borstnik:2023 , has yielded intriguing insights, with a recent surge of interest, particularly in octonions Furey:2014 ; Furey:2018a ; Furey:2018b ; Todorov:2023 ; Lasenby:2023 . While these theories demonstrate that the invariance of certain geometrical structures can give rise to symmetries related to the Standard Model (SM), their primary weakness is the ad hoc nature of the very structures involved. Consequently, their predictive power for gauge symmetries and particle representations is comparable to that of standard GUTs. Additionally, these theories typically deal with a single particle family and often overlook the quantum field theory (QFT) aspect.
@@
In this work, we depart from the historical top-bottom approach of finding mathematical frameworks that contain the SM gauge symmetries to a strict “first principles” approach. In particular, we show that the entire SM can be obtained starting from the free Dirac QFT. This is achieved by first proving that a fermionic quantum field can be constructed using solely Witt basis vectors of a $\mathcal{C}\ell_{8,0}$ Clifford algebra (via Bott periodicity), naturally representing creation-annihilation operators acting on minimal ideals. An embedding for the Dirac theory in 8-dimensional space is then obtained, unifying spinors and creation operators. The phenomenon of triality, characteristic of $\mathcal{C}\ell_{8,0}$, gives rise to two more distinct Dirac spinors representations, leading to a generalized Dirac equation with a mass matrix containing 3 families of particles per species. The minimal ideals, from which the QFT vacuum is constructed, inherit natural substructures from spinors in $\mathcal{C}\ell_{8,0}$, whose rotations have no physical consequences for the Lagrangian. This latter invariance is uniquely described by the SM group with unified gauge couplings, leading to all known particle representations plus three right handed neutrinos. The algebra of spacetime (i.e. the Dirac algebra) becomes unified in this way with the same algebra as gauge transformations, with particle multiplets related by reflections in spacetime. Ideals and Lie algebras become more than just entities associated to abstract groups, being directly connected to rotations of the quantum field vacuum which now transforms as (not just under) Clifford algebras, all the while avoiding proton decay. We call this unifying framework Spacetime Grand Unified Theory (SGUT).
@@ II Quantum Fields as Clifford Algebra Elements
However, one can actually construct fermionic creation and annihilation operators as infinite tensor products of $\mathcal{C}\ell_{8,0}$ algebras via Bott periodicity, such that $\hat{\psi}_{\alpha}(x)\in\mathcal{C}\ell_{8,0}\otimes\lim_{N\to\infty}(\mathcal{C}\ell_{8N,0})^{\otimes 3}$ (c.f. Sec. I of the Appendix). Consequently, just as relativistic invariance of quantum dynamics requires the introduction of the Clifford algebra $\mathcal{C}\ell_{1,3}$ and the Dirac equation, the quantization of its four independent solutions requires the introduction of the algebra $\mathcal{C}\ell_{8,0}$.
@@ III Relativistic Quantum Mechanics in $\mathcal{C}\ell_{8,0}$
Having singled out the importance of $\mathcal{C}\ell_{8,0}$, one might explore whether it plays some other role in the theory described by the Lagrangian in Eq. (1). Note that the latter can be equivalently written as ^$\hat{\mathcal{L}}(x)=\hat{\psi}^{\dagger}_{\alpha}(x)X_{\alpha\beta}\hat{\psi}_{\beta}(x)$ where $X_{\alpha\beta}$ are the components of a $4\times 4$ matrix $X\in\mathcal{C}\ell_{1,3}$. Since the Dirac algebra is a subalgebra of $\mathcal{C}\ell_{8,0}$, it is natural to search for $X\in\mathcal{C}\ell_{8,0}$, i.e. for an embedding of the Dirac algebra in $\mathcal{C}\ell_{8,0}$. The latter question is equivalent to finding a set of four 16-dimensional matrices $\Gamma^{\mu}_{(+)}$ and a 16-dimensional diagonal matrix $P_{(+)}$ with only four non-zero elements equal to 1, such that
@@
where $n$ and $w$ are a vector and a linear combination of $e_{A}e_{B}e_{C}e_{D}$, respectively, such that $n^{2}=1$, ${w^{2}=7+6w}$ and $[w,n]=0$. The object $w$ is commonly known as a calibration and $n$ its neutral axis. Together, the three vacua $\Omega_{(a)}$ and Witt basis vectors $\omega^{\alpha}$ generate a subspace of $C\ell_{8,0}$ via the completeness relation
@@ IV The Standard Model from $\mathcal{C}\ell_{8,0}$
Each pair $(a,A)$ gives the matrix elements appearing in the Lagrangian, so all such terms must be considered as independent quantum fields $\hat{\psi}_{\alpha,a,A}(x)$ obeying the standard Fermi anticommutation rules between themselves. For a specific particle type $(a,A)$, this amounts to taking the generalized kinetic term of Eq. (1) to be ^$\hat{\mathcal{L}}^{\textrm{Kin.}}_{a,A}(x)=\hat{\psi}^{\dagger}_{\alpha,a,A}(x)(i\gamma^{0}\gamma^{\mu})_{\alpha\beta}\partial_{\mu}\hat{\psi}_{\beta,a,A}(x)$ which, using Eq. (33) and $\textrm{Tr}[A]B=\textrm{Tr}_{A}[A\otimes B]$, can be written as
@@
and $\textrm{Tr}_{\backslash\hat{\psi}}$ denotes the trace over all subspaces which do not include quantum fields. The quantity $\omega^{\alpha\dagger}\Omega^{\dagger}_{(a)}V_{A}\in C\ell_{8,0}$ in $\hat{L}_{a,A}(x)$ plays the role of a multiplet for $\hat{\psi}_{\alpha,a,A}(x)$. In addition, since $\hat{\psi}^{\prime}_{\alpha,a,A}(x^{\prime})=S(\Lambda)_{\alpha}{}^{\beta}\hat{\psi}_{\beta,a,A}(x)$ under a Lorentz transformation, one has $\hat{L}^{\prime}_{a,A}(x^{\prime})=\mathcal{S}(\Lambda)\hat{L}_{a,A}(x)$. The spinor transformation matrix $\mathcal{S}(\Lambda)$ commutes with the pseudoscalar
@@
$V_{1}$ 1 1 00 $\nu_{eR}$
$V_{2}$ $\textbf{3}^{r}$ 1 $-2/3$ $d^{r}_{R}$
$V_{3}$ $\textbf{3}^{g}$ 1 $-2/3$ $d^{g}_{R}$
$V_{4}$ $\textbf{3}^{b}$ 1 $4/3$ $u^{b}_{R}$
$V_{5}$ 1 1 $-2$ $e_{R}$
$V_{6}$ $\textbf{3}^{r}$ 1 $4/3$ $u^{r}_{R}$
$V_{7}$ $\textbf{3}^{b}$ 1 $-2/3$ $d^{b}_{R}$
$V_{8}$ $\textbf{3}^{g}$ 1 $4/3$ $u^{g}_{R}$
$V_{9}$ $\textbf{3}^{g}$ $\textbf{2}^{\uparrow}$ $1/3$ $u^{g}_{L}$
$V_{10}$ $\textbf{3}^{b}$ $\textbf{2}^{\downarrow}$ $1/3$ $d^{b}_{L}$
$V_{11}$ $\textbf{3}^{r}$ $\textbf{2}^{\uparrow}$ $1/3$ $u^{r}_{L}$
$V_{12}$ 1 $\textbf{2}^{\downarrow}$ $-1$ $e_{L}$
$V_{13}$ $\textbf{3}^{g}$ $\textbf{2}^{\downarrow}$ $1/3$ $d^{g}_{L}$
$V_{14}$ $\textbf{3}^{b}$ $\textbf{2}^{\uparrow}$ $1/3$ $u^{b}_{L}$
$V_{15}$ $\textbf{3}^{r}$ $\textbf{2}^{\downarrow}$ $1/3$ $d^{r}_{L}$
$V_{16}$ 1 $\textbf{2}^{\uparrow}$ $-1$ $\nu_{L}$
@@ VI Acknowledgements
G.Q. thanks the support from Fundação para a Ciência e a Tecnologia, through projects CEECIND/02474/2018, 2024.04456.CERN and FCT-Mobility/1312232346/2024-25.
@@ I. Explicit construction of fermionic quantum fields using elements of $\mathcal{C}\ell_{8,0}$
For convenience, we begin by introducing the following basis $\{e_{A}\}\,(A=1,\ldots,8)$ for $\mathcal{C}\ell_{8,0}$ [^1] Due to the considerable matrix sizes involved, we will denote each set of matrices by a linear combination of the basis in order to conserve space.
@@
where the hat notation disappears since the creation operators are now finite matrices. The anticommutation relations (A35) and (A36) correspond to a Witt basis of some Clifford algebra. To identify the latter algebra, we focus on a Witt basis $\{\omega^{\alpha}\}$ for a Clifford algebra whose anticommutation relations satisfy Eq. (5). There is a well-established literature on how to build Witt bases and the properties of such objects (see Lounesto:2001 for a modern take and references therein). For the specific case of Eq. (5), the lowest dimensional representation for $\omega^{a}$ are $16\times 16$ matrices belonging to $\mathcal{C}\ell_{8,0}$ with multiple possible metric signatures, i.e. $\mathcal{C}\ell_{l,n}$ such that $l+n=8$. Without loss of generality, we will take $\omega^{a}\in\mathcal{C}\ell_{8,0}$, which can be constructed from Eq. (A34) and Eq. (6), leading to
@@
where $s_{k}=1,\ldots,4$, each $\omega^{s_{k}}$ obeys the commutation rules of Eq. (5), $I_{e}\equiv e_{1}\ldots e_{8}$ and each $\Xi^{i}$ is a $16^{N}$-dimensional matrix. It is straightforward to check that the $4N$ elements $\Xi^{i}$ satisfy the required commutation relations for a Witt basis of $\mathcal{C}\ell_{8N,0}$, i.e. $\{\Xi^{i},\Xi^{j\dagger}\}=\delta^{ij}$, $\{\Xi^{i},\Xi^{j}\}=\{\Xi^{i\dagger},\Xi^{j\dagger}\}=0$, where $i,j=1,\ldots,N$. Using the latter relations together with Eq. (5), one obtains the correspondence
@@ II. Spin(8) triality in $\mathcal{C}\ell_{8,0}$
Among the representations of the Spin($n$) group, there are three of special relevance for this work: two spinorial ones, denoted by $S_{\pm}$ and a vectorial one, denoted by $V$. For the majority of $n$, these representations don’t have the same dimensions. However, for $n=8$, all the dimensions are equal to 8 and so any two elements of different representations uniquely define a third one in the remaining representation - a phenomenon called triality. Rather than give an exposition on triality, which can get unnecessarily technical, we shall only use the results from Lounesto:2001 which will be required for this work. Note that, since all basis matrices are explicitly stated, all results presented here can be straightforwardly verified using a symbolic manipulation software such as Mathematica.
@@
$+\left(\frac{1-I_{e}}{2}\right)\bigg{]}(w-3)\{u(1+I_{e}),n\}n(w-3)^{-1}\,,$ (A51)
@@
$\left(\begin{array}[]{cccccccccccccccc}0&-c_{1,-}&-c_{1,+}&-c_{2,0}&0&-c_{2,-}&-c_{1,0}&-c_{2,+}&-c_{3,+}&-c_{4,0}&-c_{3,-}&0&-c_{4,+}&-c_{3,0}&-c_{4,-}&0\\ 0&0&0&0&0&0&0&c_{2,-}&c_{3,-}&0&0&0&0&0&0&-c_{4,-}\\ 0&-c_{2,0}&0&0&c_{1,-}&-c_{2,+}&0&0&0&0&c_{3,+}&0&0&0&c_{4,+}&0\\ 0&c_{1,+}&c_{2,-}&0&0&-c_{1,0}&c_{1,-}&0&0&0&c_{4,0}&0&0&0&c_{3,0}&0\\ 0&0&c_{1,0}&-c_{2,+}&0&0&-c_{1,+}&c_{2,0}&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&-c_{1,-}&c_{3,0}&c_{4,+}&0&c_{3,-}&c_{4,0}-c_{4,-}&c_{3,+}&0&0\\ 0&-c_{2,+}&0&0&-c_{2,-}&c_{2,0}&0&0&0&0&0&c_{3,0}&0&0&0&c_{4,0}\\ 0&c_{1,0}&0&0&0&c_{1,+}&0&0&0&-c_{4,-}&0&-c_{4,+}&0&-c_{3,-}&0&-c_{3,+}\\ 0&-c_{4,0}&0&0&c_{4,+}&0&c_{4,-}&0&0&0&-c_{1,+}&0&0&c_{2,-}&0&c_{2,+}\\ 0&c_{3,+}&0&0&-c_{3,0}-c_{3,-}&0&0&0&0&0&-c_{2,0}&0&0&0&0&-c_{1,0}\\ 0&0&0&c_{4,-}&0&0&-c_{4,+}&-c_{3,0}&-c_{1,-}&0&0&-c_{2,-}&c_{1,0}&c_{2,+}&0&0\\ 0&0&c_{4,0}+c_{4,-}&-c_{3,+}&0&0&0&0&-c_{2,0}&c_{1,+}&0&0&0&-c_{1,-}&0&0\\ 0&c_{3,0}&c_{3,-}&0&-c_{3,+}&0&0&0&0&c_{1,-}&0&c_{2,+}&0&0&-c_{1,+}&0\\ 0&-c_{4,+}&0&0&c_{4,0}+c_{4,-}&0&0&0&0&0&0&-c_{1,0}&0&0&-c_{2,0}&0\\ 0&0&0&-c_{3,-}&0&0&c_{3,+}&c_{4,0}&c_{1,0}&c_{2,+}&0&0&c_{2,-}&0&0&c_{1,-}\\ 0&0&-c_{3,0}&c_{4,+}&0&0&c_{3,-}&0&0&-c_{2,-}&0&0&-c_{2,0}&c_{1,+}&0&0\\ \end{array}\right)$ (A128)

### v2

@@ Abstract
The Standard Model of particle physics is derived from first principles starting from the free Dirac Lagrangian. All known fermionic particle species plus three right handed neutrinos are obtained from ideals of a $\mathcal{C}\ell_{8,0}$ algebra, with gauge symmetries arising as rotations of creation-annihilation operators and vacua. Triality originates both the strong force and three particle families with a mass hierarchy. Lorentz and gauge transformations are unified while avoiding the Coleman-Mandula theorem. Chirality stems unavoidably from rotations leaving the vacua invariant, with a predicted Weinberg angle of $\sin^{2}(\theta_{W})=3/8$. The theory is anomaly-free and devoid of proton decay.
@@ I Introduction
An algebraic alternative to GUTs and SUSY has been simultaneously developed over the decades, though it has received far less attention in comparison. This line of thought, typically centered on division algebras Gunaydin:1973a ; Gunaydin:1973b ; Dixon:1994 ; Leo:1996 ; Carrion:2003 ; Baez:2012 ; Baez:2011a ; Baez:2011b ; Huerta:2011 ; Huerta:2014 ; Furey:2015 ; Burdik:2017 ; Furey:2021 ; Furey:2022a ; Singh:2023 ; Furey:2022b ; Singh:2025 ; Furey:2016 ; Stoica:2018 ; Furey:2018b ; Todorov:2023 ; Lasenby:2023 ; Gording:2020 ; Manogue:2010 ; Boyle:2014 ; Boyle:2020 ; Bhatt:2022 ; Furey:2014 ; Furey:2025a ; Furey:2025b ; Todorov:2017 ; Furey:2018a ; Gillard:2019a ; Gillard:2019b or Clifford algebras Barducci:1977 ; Casalbuoni:1979a ; Casalbuoni:1979b ; Hestenes:1982 ; Cho:1995 ; Chisholm:1996 ; Lewis:1998 ; Trayling:2001 ; Trayling:2004 ; Doran:2003 ; Pavsic:2017 ; Stoica:2020 ; Borstnik:2023 ; Zenczykowski:2015 (c.f. references therein for a more complete list), has yielded intriguing insights. In particular, it demonstrates that the invariant symmetries of certain structures can give rise to symmetries related to the Standard Model (SM). For example, it is known that the algebras $\mathcal{C}\ell(7)$ and $\mathcal{C}\ell(8)$ have enough space to accommodate one and three copies, respectively, of the particle representations of the SM Chisholm:1996 ; Trayling:2004 ; Pavsic:2017 ; Gillard:2019b ; Furey:2021 . This approach, however, is limited by the ad hoc nature of the very structures involved, often leading to either an incomplete particle content or a surplus of unobserved particles, as well as a lack of testable predictions. Additional works have also focused on the problem of particle families by searching for possible connections with triality Silagadze:1995 ; YuFen:2006 ; Manogue:2010 ; Boyle:2014 ; Furey:2014 ; Todorov:2017 ; Furey:2018a ; Gillard:2019a ; Gillard:2019b ; Gording:2020 ; Boyle:2020 ; Bhatt:2022 ; Furey:2025a ; Furey:2025b , although no definite connection has been confirmed yet, partially due to the a priori non-unique reasoning for considering the phenomenon, as well as the lack of a predicted mass hierarchy. Consequently, the predictive power of these algebraic theories is still only comparable, at best, to that of standard GUTs. Furthermore, they typically overlook the quantization aspect (notable exceptions are Cho:1995 ; Zenczykowski:2015 ; Pavsic:2017 ).
@@
In this work, we depart from the historical top-bottom approach of finding mathematical frameworks that can fit the SM gauge symmetries to a strict bottom-up first-principles approach. In particular, we show that the entire SM can be obtained starting from the free Dirac quantum field theory (QFT). This is achieved by first proving that a fermionic quantum field can be constructed using solely Witt basis vectors of a $\mathcal{C}\ell_{8,0}$ Clifford algebra, naturally representing creation-annihilation operators acting on minimal ideals. An embedding for the Dirac theory in 8-dimensional space is then obtained, unifying spinors and creation operators. The phenomenon of triality, characteristic of $\mathcal{C}\ell_{8,0}$, gives rise to two more distinct Dirac spinors representations, leading to a generalized Dirac equation with a mass matrix containing 3 families of particles per species. The minimal ideals, from which the QFT vacuum is constructed, inherit natural substructures from spinors in $\mathcal{C}\ell_{8,0}$, whose rotations have no physical consequences for the Lagrangian. This latter invariance is uniquely described by the SM group with unified gauge couplings, leading to all known particle representations plus three right handed neutrinos. The algebra of spacetime (i.e. the Dirac algebra) becomes unified in this way with the same algebra as gauge transformations, with particle multiplets related by reflections in spacetime. Ideals and Lie algebras become more than just entities associated to abstract groups, being directly connected to rotations of the quantum field vacuum which now transforms as (not just under) Clifford algebras, all the while avoiding proton decay. We call this unifying framework Spacetime Grand Unified Theory (SGUT).
@@ II Quantum Fields as Clifford Algebra Elements
However, one can actually construct fermionic creation and annihilation operators as infinite tensor products of $\mathcal{C}\ell_{8,0}$ algebras via Bott periodicity, such that $\hat{\psi}_{\alpha}(x)\in\mathcal{C}\ell_{8,0}\otimes\lim_{N\to\infty}(\mathcal{C}\ell_{8N,0})^{\otimes 3}$ (c.f. Sec. I of the Appendix). Consequently, just as relativistic invariance of quantum dynamics requires the introduction of the Clifford algebra $\mathcal{C}\ell_{1,3}$ and the Dirac equation, the quantization of its four independent solutions requires the introduction of the algebra $\mathcal{C}\ell_{8,0}$. We note that in Furey:2021 it is observed that Bott periodicity might be related to the construction of multiparticle states and possibly Fock spaces. This is then used as motivation to study $\mathcal{C}\ell_{0,8}$ as previous works using the same algebra Gillard:2019a ; Gillard:2019b (modulus metric signatures), i.e. by searching for structures whose invariant symmetries might lead to the SM, thus sharing the same setbacks as the remaining state-of-the-art. This latter approach shares no similarities with the one developed in this work, other than the focus on a $\mathcal{C}\ell(8)$ algebra (or Clifford algebras in general) as the setting to develop a theory of fundamental forces. In particular, the use of $\mathcal{C}\ell_{8,0}$ here is not postulated, emerging as a necessary consequence of second quantization, as shown by the novel proofs in Sec. I of the Appendix (which happen to use the Bott periodicity theorem). The works most closely aligned to this approach are Zenczykowski:2015 , where (first) quantization is connected to $\mathcal{C}\ell_{6,0}$ and Cho:1995 ; Pavsic:2017 , which considered explicit representations of fermionic quantum fields, living in the algebras $\mathcal{C}\ell_{1,3}\otimes\mathcal{C}\ell_{\infty}$ and $\mathcal{C}\ell_{8,0}\otimes\mathcal{C}\ell_{\infty}$, respectively. The proposal that the SM symmetries and particle representations might be embedded in $\mathcal{C}\ell(8)$ goes at least as far back as 2010, in the concluding remarks of Pavsic:2010 .
@@ III Relativistic Quantum Mechanics in $\mathcal{C}\ell_{8,0}$
Having singled out the importance of $\mathcal{C}\ell_{8,0}$, one might explore whether it plays some other role in the theory described by the Lagrangian in Eq. (1). Note that the latter can be equivalently written as ^$\hat{\mathcal{L}}(x)=\hat{\psi}^{\dagger}_{\alpha}(x)X_{\alpha\beta}\hat{\psi}_{\beta}(x)$ where $X_{\alpha\beta}$ are the components of a $\times 4$ 4 matrix $X\in\mathcal{C}\ell_{1,3}$. Since the Dirac algebra is a subalgebra of $\mathcal{C}\ell_{8,0}$, it is natural to search for $X\in\mathcal{C}\ell_{8,0}$, i.e. for an embedding of the Dirac algebra in $\mathcal{C}\ell_{8,0}$. The latter question is equivalent to finding a set of four 16-dimensional matrices $\Gamma^{\mu}_{(+)}$ and a 16-dimensional diagonal matrix $P_{(+)}$ with only four non-zero elements equal to 1, such that
@@
[…]
@@ IV The Standard Model from $\mathcal{C}\ell_{8,0}$
[…]
@@
[…]
@@
$V_{1}$ 1 1 0 $\nu_{eR}$
$V_{2}$ $\textbf{3}^{r}$ 1 2/3-2/3- 2 / 3 $d^{r}_{R}$
$V_{3}$ $\textbf{3}^{g}$ 1 2/3-2/3- 2 / 3 $d^{g}_{R}$
$V_{4}$ $\textbf{3}^{b}$ 1 4/34/34 / 3 $u^{b}_{R}$
$V_{5}$ 1 1 2-2- 2 $e_{R}$
$V_{6}$ $\textbf{3}^{r}$ 1 4/34/34 / 3 $u^{r}_{R}$
$V_{7}$ $\textbf{3}^{b}$ 1 2/3-2/3- 2 / 3 $d^{b}_{R}$
$V_{8}$ $\textbf{3}^{g}$ 1 4/34/34 / 3 $u^{g}_{R}$
$V_{9}$ $\textbf{3}^{g}$ $\textbf{2}^{\uparrow}$ 1/31/31 / 3 $u^{g}_{L}$
$V_{10}$ $\textbf{3}^{b}$ $\textbf{2}^{\downarrow}$ 1/31/31 / 3 $d^{b}_{L}$
$V_{11}$ $\textbf{3}^{r}$ $\textbf{2}^{\uparrow}$ 1/31/31 / 3 $u^{r}_{L}$
$V_{12}$ 1 $\textbf{2}^{\downarrow}$ 1-1- 1 $e_{L}$
$V_{13}$ $\textbf{3}^{g}$ $\textbf{2}^{\downarrow}$ 1/31/31 / 3 $d^{g}_{L}$
$V_{14}$ $\textbf{3}^{b}$ $\textbf{2}^{\uparrow}$ 1/31/31 / 3 $u^{b}_{L}$
$V_{15}$ $\textbf{3}^{r}$ $\textbf{2}^{\downarrow}$ 1/31/31 / 3 $d^{r}_{L}$
$V_{16}$ 1 $\textbf{2}^{\uparrow}$ 1-1- 1 $\nu_{L}$
@@ VI Acknowledgements
I thank the support from Fundação para a Ciência e a Tecnologia, through projects CEECIND/02474/2018, 2024.04456.CERN and FCT-Mobility/1312232346/2024-25. I also thank everyone who gave useful feedback in an earlier version of the paper, leading to a more complete state-of-the-art review. In particular, I thank Niels Gresnigt for pointing out to me reference Zenczykowski:2015 . Lastly, I thank Filipe Joaquim for fruitful discussions regarding GUTs.
@@ I. Explicit construction of fermionic quantum fields using elements of $\mathcal{C}\ell_{8,0}$
In this sction we prove for the first time that fermionic quantum fields can be constructed as infinite tensor products of $\mathcal{C}\ell_{8,0}$ algebras. For convenience, we begin by introducing the following basis $\{e_{A}\}\,(A=1,\ldots,8)$ for $\mathcal{C}\ell_{8,0}$ [^1] Due to the considerable matrix sizes involved, we will denote each set of matrices by a linear combination of the basis in order to conserve space.
@@
where the hat notation disappears since the creation operators are now finite matrices. The anticommutation relations (A35) and (A36) correspond to a Witt basis of some Clifford algebra. To identify the latter algebra, we focus on a Witt basis $\{\omega^{\alpha}\}$ for a Clifford algebra whose anticommutation relations satisfy Eq. (5). There is a well-established literature on how to build Witt bases and the properties of such objects (see Lounesto:2001 for a modern take and references therein). For the specific case of Eq. (5), the lowest dimensional representation for $\omega^{a}$ are $\times 16$ 16 matrices belonging to $\mathcal{C}\ell_{8,0}$ with multiple possible metric signatures, i.e. $\mathcal{C}\ell_{l,n}$ such that $l+n=8$. Without loss of generality, we will take $\omega^{a}\in\mathcal{C}\ell_{8,0}$, which can be constructed from Eq. (A34) and Eq. (6), leading to
@@
[…]
@@ II. Spin(8) triality in $\mathcal{C}\ell_{8,0}$
[…]
@@
$\hskip-22.76219pt+\left(\frac{1-I_{e}}{2}\right)\bigg{]}(w-3)\{u(1+I_{e}),n\}n(w-3)^{-1}\,,$ (A51)
@@
$\hskip-25.60747pt\left(\begin{array}[]{cccccccccccccccc}0&-c_{1,-}&-c_{1,+}&-c_{2,0}&0&-c_{2,-}&-c_{1,0}&-c_{2,+}&-c_{3,+}&-c_{4,0}&-c_{3,-}&0&-c_{4,+}&-c_{3,0}&-c_{4,-}&0\\ 0&0&0&0&0&0&0&c_{2,-}&c_{3,-}&0&0&0&0&0&0&-c_{4,-}\\ 0&-c_{2,0}&0&0&c_{1,-}&-c_{2,+}&0&0&0&0&c_{3,+}&0&0&0&c_{4,+}&0\\ 0&c_{1,+}&c_{2,-}&0&0&-c_{1,0}&c_{1,-}&0&0&0&c_{4,0}&0&0&0&c_{3,0}&0\\ 0&0&c_{1,0}&-c_{2,+}&0&0&-c_{1,+}&c_{2,0}&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&-c_{1,-}&c_{3,0}&c_{4,+}&0&c_{3,-}&c_{4,0}-c_{4,-}&c_{3,+}&0&0\\ 0&-c_{2,+}&0&0&-c_{2,-}&c_{2,0}&0&0&0&0&0&c_{3,0}&0&0&0&c_{4,0}\\ 0&c_{1,0}&0&0&0&c_{1,+}&0&0&0&-c_{4,-}&0&-c_{4,+}&0&-c_{3,-}&0&-c_{3,+}\\ 0&-c_{4,0}&0&0&c_{4,+}&0&c_{4,-}&0&0&0&-c_{1,+}&0&0&c_{2,-}&0&c_{2,+}\\ 0&c_{3,+}&0&0&-c_{3,0}-c_{3,-}&0&0&0&0&0&-c_{2,0}&0&0&0&0&-c_{1,0}\\ 0&0&0&c_{4,-}&0&0&-c_{4,+}&-c_{3,0}&-c_{1,-}&0&0&-c_{2,-}&c_{1,0}&c_{2,+}&0&0\\ 0&0&c_{4,0}+c_{4,-}&-c_{3,+}&0&0&0&0&-c_{2,0}&c_{1,+}&0&0&0&-c_{1,-}&0&0\\ 0&c_{3,0}&c_{3,-}&0&-c_{3,+}&0&0&0&0&c_{1,-}&0&c_{2,+}&0&0&-c_{1,+}&0\\ 0&-c_{4,+}&0&0&c_{4,0}+c_{4,-}&0&0&0&0&0&0&-c_{1,0}&0&0&-c_{2,0}&0\\ 0&0&0&-c_{3,-}&0&0&c_{3,+}&c_{4,0}&c_{1,0}&c_{2,+}&0&0&c_{2,-}&0&0&c_{1,-}\\ 0&0&-c_{3,0}&c_{4,+}&0&0&c_{3,-}&0&0&-c_{2,-}&0&0&-c_{2,0}&c_{1,+}&0&0\\ \end{array}\right)$ (A128)
As far as the author is aware, all results from Eq. (A60) to Eq. (II. Spin(8) triality in $\mathcal{C}\ell_{8,0}$) are completely novel results in the Clifford algebra literature.